# A Comprehensive Overview of Large Language Models

### 20 Feb 2024 v8

@@ III-H2 Training Strategies
Training models at a huge scale require tricks to reduce training costs, avoid loss divergence and achieve better performance. We summarize and discuss some of these key tricks used in different LLMs.
@@
Mixed Precision: is a famous method for LLMs to reduce memory usage and improve training efficiency. In mixed precision, forward and backward passes are performed in FP16 format whereas optimizer states and master weights are kept in FP32 format [115]. A drawback associated with this format change is training instability due to a smaller value range resulting in loss spikes [33]. An alternative to FP16 is BF16 which has a comparatively larger range and performs some precision-sensitive operations like gradient accumulation and softmax in FP32 [13]. BF16 has better performance and training stability but uses more memory and is supported on specific hardware, for example, A100 GPUs. Therefore, its adoption in LLMs is limited.
@@
Training Instability: is a common issue in LLMs where loss divergence or spiking is observed multiple times during training. This happens in the presence of gradient clipping [15]. To mitigate this problem, many approaches suggest restarting training from an earlier checkpoint [15, 33, 116], skipping 200-500 earlier data batches at the point of divergence in [15] and re-shuffling batches in [116]. The embedding layer gradient shrink proves to further stabilize the training as its gradient norm is significantly larger than the other layers [33]. Another suggestion to improve training stability for larger models is not to use biases in dense and norm layers as in [15].
@@
Weight Initialization: plays a significant role in model convergence and training stability. GPT-NeoX [113] initializes feed-forward layers before residuals with $\frac{2}{L\sqrt{d}}$ as in [142] and other layers with small initialization scheme [287]. This avoids activations growing exponentially with the increasing depth. MT-NLG [112] found higher variance for weight initialization leads to unstable training, hence validating small initialization scheme [287]. Various models perform random weight initialization which can cause bad initialization, Galactica [137] suggests a longer warmup to negate the effect.
@@
Learning Rate: is important for stable training. It is suggested to use a lower value [13, 15, 122] with warmup and decay (cosine or linear). Usually, the learning rate is within the range $1e^{-4}$ to $8e^{-4}$. Moreover, MT-NLG (530B) [112] and GPT-NeoX (20B) [113] suggest interpolating learning rates based on the model size using the GPT-3 [6] models ranging between 13B and 175B. This avoids tuning the learning rate hyperparameter.
@@
Training Parallelism: 3D parallelism, a combination of data, pipeline and tensor parallelism, is the most utilized training parallelism approach in LLMs [33, 15, 14, 13, 112, 110, 107]. In addition to the 3D parallelism, BLOOM [13] uses zero optimizer [37] to shard optimizer states. PanGu-$\alpha$ [103] and PanGu-$\Sigma$ [128] go beyond the 3D parallelism and apply 5D parallelism which additionally contains optimizer parallelism and rematerialization.
@@
Mode Switching: adds task-related tokens at the beginning of the text during training. These tokens refer to the natural language understanding and natural language generation tasks which are shown to improve the downstream task performance in [123, 122, 120]. During fine-tuning and inference, tokens are appended based on the downstream tasks.
@@
III-H3 Supervised Models vs Generalized Models
@@
III-H4 Zero-Shot vs Few-Shot
LLMs perform well in zero-shot and few-shot settings. But the performance difference between zero-shot and few-shot is large for pre-trained models [6, 15], naming LLMs as meta-learners [6]. LLMs zero-shot evaluations underperform unsupervised methods in neural machine translation [6]. The literature shows pre-training is not enough for good zero-shot performance [15, 16]. To improve the zero-shot performance the literature suggests using instruction fine-tuning that improves the zero-shot performance significantly and outperforms baselines. Instruction fine-tuning has also been shown to improve zero-shot generalization to unseen tasks. Another model Flan-PaLM [16] unlocks zero-shot reasoning with CoT training.
III-H5 Encoder vs Decoder vs Encoder-Decoder
Traditionally, these architectures perform well for different tasks, for example, encoder-only for NLU tasks, decoder-only for NLG, and encoder-decoder for sequence2sequence modeling. Encoder-only models are famous for smaller models such as Bert [7], RoBERTa [288], etc, whereas LLMs are either decoder-only [6, 113, 13] or encoder-decoder [10, 11, 120]. While decoder-only models are good at NLG tasks, various LLMs, PaLM [15], OPT [14], GPT-3 [6], BLOOM [13], LLaMA [145], are decoder-only models with significant performance gains on both NLU and NLG tasks. In contradiction to this, T5 [10] and UL2 [123] identify encoder-decoder models out-performing decoder-only models. In another study, PaLM [15] finds increasing the size of decoder-only models can reduce the performance gap between decoder-only and encoder-decoder architectures.
@@ III-H5 Encoder vs Decoder vs Encoder-Decoder
TABLE III: Summary of pre-trained LLMs (>10B). Only the LLMs discussed individually in the previous sections are summarized. “Data/Tokens” is the model’s pre-training data, which is either the number of tokens or data size. “Data Cleaning” indicates whether the data cleaning is performed or not. This includes heuristics (Heur), deduplication (Dedup), quality filtering (QF), and privacy filtering (PF), “Cost” is the calculated training cost obtained by multiplying the GPUs/TPUs hourly rate with the number of GPUs and the training time. The actual cost may vary due to many reasons such as using in-house GPUs or getting a discounted rate, re-training, number of employees working on the problem, etc. “Training Parallelism” indicates distributed training using data parallelism (D), tensor parallelism (T), pipeline parallelism (P), model parallelism (M), optimizer parallelism (OP), and rematerialization (R), where for “Library” column, “DS” is a short form for Deep Speed. In column “Commercial Use”, we assumed a model is for non-commercial purposes if its license is unavailable.
Models Publication Venue License Type Model Creators Purpose No. of Params   Commercial Use Steps Trained Data/ Tokens Data Cleaning No. of Processing Units Processing Unit Type Training Time Calculated Train. Cost Training Parallelism Library
@@
ERNIE 3.0 [105] arXiv'21 - Baidu General 10B $\times$ $*{}^{*}$ 375B Heur+Dedup 384 V100 - - $*{}^{*}$ PaddlePaddle
@@
Yuan 1.0 [110] arXiv'21 Apache-2.0 - General 245B $\checkmark$ $*{}^{*}$ 180B Heur+Clf+Dedup 2128 GPU - - D+T+P -
@@
GLaM [116] ICML'22 - Google General 1.2T $\times$ $*{}^{*}$ 600B Clf 1024 TPU v4 - - M GSPMD
@@
TABLE IV: Summary of instruction tuned LLMs (>10B). All abbreviations are the same as Table III. Entries in “Data/Tokens” starting with “S-” represents the number of training samples.
Models Publication Venue License Type Model Creators Purpose No. of Params Commercial Use Pre-trained Models Steps Trained Data/ Tokens No. of Processing Units Processing Unit Type Train. Time Calculated Train. Cost Train. Parallelism Library
@@
IV Model Configurations
We provide different statistics of pre-trained and instruction-tuned models in this section. This includes information such as publication venue, license type, model creators, steps trained, parallelism, etc in Table III and Table IV. Architecture details of pre-trained LLMs are available in Table V. Providing these details for instruction-tuned models is unnecessary because it fine-tunes pre-trained models for instruction datasets. Hence, architectural details are the same as the baselines. Moreover, optimization settings for various LLMs are available in Table VI and Table VII. We do not include details on precision, warmup, and weight decay in Table VII. Neither of these details are important as others to mention for instruction-tuned models nor provided by the papers.
TABLE V: Architecture details of LLMs. Here, “PE” is the positional embedding, “nL” is the number of layers, “nH” is the number of attention heads, “HS” is the size of hidden states.
Models Type
@@ IV Model Configurations
Jurassic-1 (178B) Causal-Dec Next Token Standard 256k $*{}^{*}$ Pre-Layer Learned GeLU $\checkmark$ 76 96 13824
@@
Sparrow (70B) Causal-Dec Pref.&Rule RM - 32k SentencePiece-NFKC Pre-RMS Relative GeLU $\checkmark$ $*{}^{*}$ 64 8192
@@
TABLE VI: Summary of optimization settings used for pre-trained LLMs. The values for weight decay, gradient clipping, and dropout are 0.1, 1.0, and 0.1, respectively, for most of the LLMs.
Sequence LR Optimizers Precision Weight Grad
Models Batch Size Length LR Warmup Decay AdaFactor Adam AdamW FP16 BF16 Mixed Decay Clip Dropout
@@
$\checkmark$
- - $\checkmark$
GPT3 (175B) 32K - 6e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ -
mT5 (13B) 1024 1024 0.01 - inverse square root
$\checkmark$
- - $\checkmark$
PanGu-$\alpha$ (200B) - 1024 2e-5 - -
$\checkmark$
CPM-2 (198B) 1024 1024 0.001 - -
$\checkmark$
- - $\checkmark$
Codex (12B) - - 6e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ - -
ERNIE 3.0 (12B) 6144 512 1e-4 $\checkmark$ linear
$\checkmark$
$\checkmark$ - -
Jurassic-1 (178B) 3.2M 2048 6e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ -
HyperCLOVA (82B) 1024 - 6e-5 - cosine
$\checkmark$
$\checkmark$ - -
Yuan 1.0 (245B) $<$10M 2048 1.6e-4 $\checkmark$ cosine decay to 10%
$\checkmark$
$\checkmark$ - -
Gopher (280B) 3M 2048 4e-5 $\checkmark$ cosine decay to 10%
$\checkmark$
$\checkmark$
- $\checkmark$ -
ERNIE 3.0 Titan (260B) - 512 1e-4 $\checkmark$ linear
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ -
GPT-NeoX-20B 1538 2048 0.97e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ $\times$
OPT (175B) 2M 2048 1.2e-4 - linear
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ $\checkmark$
BLOOM (176B) 2048 2048 6e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ $\times$
Galactica (120B) 2M 2048 7e-6 $\checkmark$ linear decay to 10%
$\checkmark$
$\checkmark$ $\checkmark$ $\checkmark$
GLaM (1.2T) 1M 1024 0.01 - inverse square root
$\checkmark$
FP32 + $\checkmark$ - $\checkmark$ $\times$
LaMDA (137B) 256K - - - -
MT-NLG (530B) 1920 2048 5e-5 $\checkmark$ cosine decay to 10%
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ -
AlphaCode (41B) 2048 1536+768 1e-4 $\checkmark$ cosine decay to 10%
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ -
Chinchilla (70B) 1.5M 2048 1e-4 $\checkmark$ cosine decay to 10%
$\checkmark$
$\checkmark$
PaLM (540B) 2048 2048 0.01 - inverse square root
$\checkmark$
$\checkmark$ $\checkmark$ $\times$
AlexaTM (20B) 2M 1024 1e-4 - linear decay to 5%
$\checkmark$
$\checkmark$
$\checkmark$ - $\checkmark$
U-PaLM (540B) 32 2048 1e-4 - cosine
$\checkmark$
UL2 (20B) 1024 1024 - - inverse square root
$\times$ - -
GLM (130B) 4224 2048 8e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ $\checkmark$
CodeGen (16B) 2M 2048 5e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$ $\checkmark$ -
LLaMA (65B) 4M Tokens 2048 1.5e-4 $\checkmark$ cosine decay to 10%
$\checkmark$
$\checkmark$ $\checkmark$ -
PanGu-$\Sigma$ (1.085T) 512 1024 2e-5 $\checkmark$ -
$\checkmark$
$\checkmark$
BloombergGPT (50B) 2048 2048 6e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ $\times$
Xuan Yuan 2.0 (176B) 2048 2048 6e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ -
CodeT5+ (16B) 2048 1024 2e-4 - linear
$\checkmark$
$\checkmark$
$\checkmark$ - -
StarCoder (15.5B) 512 8k 3e-4 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ - -
LLaMA-2 (70B) 4M Tokens 4k 1.5e-4 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ -
TABLE VII: Summary of optimization settings used for instruction-tuned LLMs. Values for gradient clipping and dropout are the same as the pre-trained models, while no model uses weight decay for instruction tuning.
Sequence Optimizers Grad
Models Batch Size Length LR Warmup LR_Decay AdaFactor Adam AdamW Clip Dropout

### 09 Apr 2024 v9

@@ 3.8.2 Training Strategies
Training models at a huge scale require tricks to reduce training costs, avoid loss divergence, and achieve better performance. We summarize and discuss some of these key tricks used in different LLMs.
@@
Mixed Precision: It is a famous method for LLMs to reduce memory usage and improve training efficiency. In mixed precision, forward and backward passes are performed in FP16 format whereas optimizer states and master weights are kept in FP32 format [120]. A drawback associated with this format change is training instability due to a smaller value range resulting in loss spikes [33]. An alternative to FP16 is BF16 which has a comparatively larger range and performs precision-sensitive operations like gradient accumulation and softmax in FP32 [13]. BF16 has better performance and training stability but uses more memory and is supported on specific hardware, for example, A100 GPUs. Therefore, its adoption in LLMs is limited.
@@
Training Instability: Loss divergence or spiking is a common issue in LLMs that occurs multiple times during training. This happens in the presence of gradient clipping [15]. To mitigate this problem, many approaches suggest restarting training from an earlier checkpoint [15, 33, 91], skipping 200-500 earlier data batches at the point of divergence in [15] and re-shuffling batches in [91]. The embedding layer gradient shrink proves to further stabilize the training as its gradient norm is significantly larger than the other layers [33]. Another suggestion to improve training stability for larger models is not to use biases in dense and norm layers as in [15].
@@
Weight Initialization: It plays a significant role in model convergence and training stability. GPT-NeoX [118] initializes feed-forward layers before residuals with $\frac{2}{L\sqrt{d}}$ as in [143] and other layers with the small initialization scheme [288]. This avoids activations growing exponentially with increasing depth. MT-NLG [117] found higher variance for weight initialization leads to unstable training, hence validating small initialization scheme [288]. Various models perform random weight initialization which can cause bad initialization, Galactica [138] suggests a longer warmup to negate the effect.
@@
Learning Rate: A suitable learning rate is important for stable training. It is suggested to use a lower value [13, 15, 124] with warmup and decay (cosine or linear). Usually, the learning rate is within the range $1e^{-4}$ to $8e^{-4}$. Moreover, MT-NLG (530B) [117] and GPT-NeoX (20B) [118] suggest interpolating learning rates based on the model size using the GPT-3 [6] models ranging between 13B and 175B. This avoids tuning the learning rate hyperparameter.
@@
Training Parallelism: 3D parallelism, a combination of data, pipeline, and tensor parallelism, is the most utilized training parallelism approach in LLMs [33, 15, 14, 13, 117, 115, 112]. In addition to 3D parallelism, BLOOM [13] uses a zero optimizer [37] to shard optimizer states. PanGu-$\alpha$ [108] and PanGu-$\Sigma$ [92] go beyond 3D parallelism and apply 5D parallelism which additionally contains optimizer parallelism and rematerialization.
@@
Mode Switching: It adds task-related tokens at the beginning of the text during training. These tokens refer to the natural language understanding and natural language generation tasks which are shown to improve downstream task performance in [125, 124, 122]. During fine-tuning and inference, tokens are appended based on the downstream tasks.
@@
3.8.3 Supervised Models vs Generalized Models
@@
3.8.4 Zero-Shot vs Few-Shot
LLMs perform well in zero-shot and few-shot settings. But the performance difference between zero-shot and few-shot is large for pre-trained models [6, 15], naming LLMs as meta-learners [6]. LLMs zero-shot evaluations underperform unsupervised methods in neural machine translation [6]. The literature shows pre-training is not enough for good zero-shot performance [15, 16]. To improve the zero-shot performance the literature suggests using instruction fine-tuning that improves the zero-shot performance significantly and outperforms baselines. Instruction fine-tuning has also been shown to improve zero-shot generalization to unseen tasks. Another model, Flan-PaLM [16], unlocks zero-shot reasoning with CoT training.
3.8.5 Encoder vs Decoder vs Encoder-Decoder
Traditionally, these architectures perform well for different tasks, for example, encoder-only for NLU tasks, decoder-only for NLG, and encoder-decoder for sequence2sequence modeling. Encoder-only models are famous for smaller models such as Bert [7], RoBERTa [289], etc., whereas LLMs are either decoder-only [6, 118, 13] or encoder-decoder [10, 11, 122]. While decoder-only models are good at NLG tasks, various LLMs, PaLM [15], OPT [14], GPT-3 [6], BLOOM [13], LLaMA [146], are decoder-only models with significant performance gains on both NLU and NLG tasks. In contradiction to this, T5 [10] and UL2 [125] identify encoder-decoder models out-performing decoder-only models. In another study, PaLM [15] finds increasing the size of decoder-only models can reduce the performance gap between decoder-only and encoder-decoder architectures.
@@ 3.8.5 Encoder vs Decoder vs Encoder-Decoder
Table 3: Summary of pre-trained LLMs (>10B). Only the LLMs discussed individually in the previous sections are summarized. “Data/Tokens” is the model’s pre-training data, which is either the number of tokens or data size. “Data Cleaning” indicates whether data cleaning is performed or not. This includes heuristics (Heur), deduplication (Dedup), quality filtering (QF), and privacy filtering (PF), “Cost” is the calculated training cost obtained by multiplying the GPUs/TPUs hourly rate with the number of GPUs and the training time. The actual cost may vary due to many reasons such as using in-house GPUs or getting a discounted rate, re-training, number of employees working on the problem, etc. “Training Parallelism” indicates distributed training using data parallelism (D), tensor parallelism (T), pipeline parallelism (P), model parallelism (M), optimizer parallelism (OP), and rematerialization (R), where for “Library” column, “DS” is a short form for Deep Speed. In column “Commercial Use”, we assumed a model is for non-commercial purposes if its license is unavailable.
[…]
@@
ERNIE 3.0 [110] arXiv'21 - Baidu General 10B $\times$ 120k 375B Heur+Dedup 384 V100 - - M PaddlePaddle
@@
Yuan 1.0 [115] arXiv'21 Apache-2.0 - General 245B $\checkmark$ 26k 180B Heur+Clf+Dedup 2128 GPU - - D+T+P -
@@
GLaM [91] ICML'22 - Google General 1.2T $\times$ 600k 600B Clf 1024 TPU v4 - - M GSPMD
@@
Table 4: Summary of instruction tuned LLMs (>10B). All abbreviations are the same as Table 3. Entries in “Data/Tokens” starting with “S-” represents the number of training samples.
[…]
@@
4 Model Configurations
We provide different statistics of pre-trained and instruction-tuned models in this section. This includes information such as publication venue, license type, model creators, steps trained, parallelism, etc in Table 3 and Table 4. Architecture details of pre-trained LLMs are available in Table 5. Providing these details for instruction-tuned models is unnecessary because it fine-tunes pre-trained models for instruction datasets. Hence, architectural details are the same as the baselines. Moreover, optimization settings for various LLMs are available in Table 6 and Table 7. We do not include details on precision, warmup, and weight decay in Table 7. These details are not as important as others to mention for instruction-tuned models, and are not provided by the papers.
Table 5: Architecture details of LLMs. Here, “PE” is the positional embedding, “nL” is the number of layers, “nH” is the number of attention heads, “HS” is the size of hidden states.
[…]
@@ 4 Model Configurations
Jurassic-1 (178B) Causal-Dec Next Token Standard 256k SentencePiece Pre-Layer Learned GeLU $\checkmark$ 76 96 13824
@@
Sparrow (70B) Causal-Dec Pref.&Rule RM - 32k SentencePiece-NFKC Pre-RMS Relative GeLU $\checkmark$ 16 64 8192
@@
Table 6: Summary of optimization settings used for pre-trained LLMs. The values for weight decay, gradient clipping, and dropout are 0.1, 1.0, and 0.1, respectively, for most of the LLMs.
@@
[…]
T5 (11B) 211 512 0.01 $\times$ inverse square root $\checkmark$ - - - - - $\checkmark$
GPT3 (175B) 32K - 6e-5 $\checkmark$ cosine $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ -
mT5 (13B) 1024 1024 0.01 - inverse square root $\checkmark$ - - - - - $\checkmark$
PanGu-$\alpha$ (200B) - 1024 2e-5 - - - - - - $\checkmark$ - - - -
CPM-2 (198B) 1024 1024 0.001 - - $\checkmark$ - - - - - $\checkmark$
Codex (12B) - - 6e-5 $\checkmark$ cosine $\checkmark$ $\checkmark$ $\checkmark$ - -
ERNIE 3.0 (12B) 6144 512 1e-4 $\checkmark$ linear $\checkmark$ - - - $\checkmark$ - -
Jurassic-1 (178B) 3.2M 2048 6e-5 $\checkmark$ cosine $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ -
HyperCLOVA (82B) 1024 - 6e-5 - cosine $\checkmark$ - - - $\checkmark$ - -
Yuan 1.0 (245B) <<10M 2048 1.6e-4 $\checkmark$ cosine decay to 10% $\checkmark$ - - - $\checkmark$ - -
Gopher (280B) 3M 2048 4e-5 $\checkmark$ cosine decay to 10% $\checkmark$ $\checkmark$ - $\checkmark$ -
ERNIE 3.0 Titan (260B) - 512 1e-4 $\checkmark$ linear $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ -
GPT-NeoX-20B 1538 2048 0.97e-5 $\checkmark$ cosine $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
OPT (175B) 2M 2048 1.2e-4 - linear $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
BLOOM (176B) 2048 2048 6e-5 $\checkmark$ cosine $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
Galactica (120B) 2M 2048 7e-6 $\checkmark$ linear decay to 10% $\checkmark$ - - - $\checkmark$ $\checkmark$ $\checkmark$
GLaM (1.2T) 1M 1024 0.01 - inverse square root $\checkmark$ FP32 + $\checkmark$ - $\checkmark$ $\times$
LaMDA (137B) 256K - - - - - - - - - - - - -
MT-NLG (530B) 1920 2048 5e-5 $\checkmark$ cosine decay to 10% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ -
AlphaCode (41B) 2048 1536+768 1e-4 $\checkmark$ cosine decay to 10% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ -
Chinchilla (70B) 1.5M 2048 1e-4 $\checkmark$ cosine decay to 10% $\checkmark$ $\checkmark$ - - -
PaLM (540B) 2048 2048 0.01 - inverse square root $\checkmark$ - - - $\checkmark$ $\checkmark$ $\times$
AlexaTM (20B) 2M 1024 1e-4 - linear decay to 5% $\checkmark$ $\checkmark$ $\checkmark$ - $\checkmark$
U-PaLM (540B) 32 2048 1e-4 - cosine $\checkmark$ - - - - - -
UL2 (20B) 1024 1024 - - inverse square root - - - - - - $\times$ - -
GLM (130B) 4224 2048 8e-5 $\checkmark$ cosine $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CodeGen (16B) 2M 2048 5e-5 $\checkmark$ cosine $\checkmark$ - - - $\checkmark$ $\checkmark$ -
LLaMA (65B) 4M Tokens 2048 1.5e-4 $\checkmark$ cosine decay to 10% $\checkmark$ - - - $\checkmark$ $\checkmark$ -
PanGu-$\Sigma$ (1.085T) 512 1024 2e-5 $\checkmark$ - $\checkmark$ $\checkmark$ - - -
BloombergGPT (50B) 2048 2048 6e-5 $\checkmark$ cosine $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
Xuan Yuan 2.0 (176B) 2048 2048 6e-5 $\checkmark$ cosine $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ -
CodeT5+ (16B) 2048 1024 2e-4 - linear $\checkmark$ $\checkmark$ $\checkmark$ - -
StarCoder (15.5B) 512 8k 3e-4 $\checkmark$ cosine $\checkmark$ $\checkmark$ $\checkmark$ - -
LLaMA-2 (70B) 4M Tokens 4k 1.5e-4 $\checkmark$ cosine $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ -
Table 7: Summary of optimization settings used for instruction-tuned LLMs. Values for gradient clipping and dropout are the same as the pre-trained models, while no model uses weight decay for instruction tuning.
[…]